OpenStudy (anonymous):

f(x) = 3x + 6, g(x) = 2x2 Find (fg)(x).

OpenStudy (waleed_imtiaz):

fg(x) means composition.. So according to me, It will be..... fg(x)=f[g(x)]=f[2x^2]=3(2x^2)+6=12x^2+6 will be the asnwer.... what U got ?

hartnn (hartnn):

3*2x^2 is 12x^2 ??? great :P

OpenStudy (anonymous):

Answer: f(2x^2)=6x^2+6

OpenStudy (waleed_imtiaz):

ya.6x^2+6

OpenStudy (anonymous):

yeah

OpenStudy (waleed_imtiaz):

its that when I solve that 2 has also the power of square..... got it ? :P

OpenStudy (anonymous):

these are my inserts tho...

OpenStudy (anonymous):

a 2x^2 + 3x + 6 b6x + 12 c 6x^2 + 12x d6x^3 + 12x^2

OpenStudy (anonymous):

its A